OpenStudy (anonymous):

Solve the equation for x and round your answer to the nearest hundreth. 5 = ln(x + 1)

OpenStudy (mathmale):

Think of 5 and ln (x+1) as exponents of the base ' e '. What is \[e^5=e ^{\ln (x+1)}?\]

OpenStudy (sidsiddhartha):

use A=ln x x=e^A

OpenStudy (mathmale):

Note that "ln" means "logarithm to the base ' e '"